OpenStudy (justin12):

A party rental company has chairs and tables for rent. The total cost to rent 6 chairs and 2 tables is $20. The total cost to rent 3 chairs and 5 tables is $35. What is the cost to rent each chair and each table?

OpenStudy (horse_rider_133):

what are the answer choices

OpenStudy (justin12):

there's none

OpenStudy (horse_rider_133):

oh ok

OpenStudy (justin12):

But I got x = 5/4 and y = 25/4 idk if thats right

OpenStudy (horse_rider_133):

let me see

OpenStudy (horse_rider_133):

here is the answer 3c+5t = 35 6c +2t = 20 Multiply eq 1 by 2 6c + 10 t = 70 Subtract eq 2 8t = 50 t = 6.25 6c + 2*6.25 = 20 C = 1.25 @justin12